OpenStudy (anonymous):

Part A: A = 9 over 2 (x + 23) solve for x.

OpenStudy (anonymous):

\[A=\frac{ 9 }{ 2(x+23) }\] like this friend?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

There's a few ways you can do this, I guess you can distribute if you want, either way will be equally length I guess. So distribute 2 over the x and 23 which will give you \[2(x+23) \implies 2x + 46\] \[A = \frac{ 9 }{ 2x+46 }\] cross multiply \[2x+46 = \frac{ 9 }{ A }\] \[2x = \frac{ 9 }{ A }-46\] we cancel 46 on the left side by subtracting, so now just divide the whole equation by 2 for your final answer.

OpenStudy (catlover5925):

i got X = 2/9 a - 23

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

There's a problem here, you assumed they gave you the right answer rather than looking through the steps I gave you, and with this I know you don't care so there is no need for me to tell you whether the answer you think is right or not.

OpenStudy (anonymous):

its right i look at the steps i did the problem i just didnt know if the got the right anwser or no that it

OpenStudy (isaiah.feynman):

Wow the professor can't take it anymore!

OpenStudy (anonymous):

Not quite, the answer Cat gave you was wrong. You can clearly see x=9/2A-23. Lol @Isaiah.Feynman :p